OpenStudy (anonymous):

Think of a 5 digit number and call it N. When N has a 7 placed in front of it, the number is 5 times as large as when N has a 7 placed after it. What number is N?

OpenStudy (anonymous):

You equation is \[ 7\cdot 10^5+N=5\cdot(10\cdot N+7) \]

OpenStudy (anonymous):

So your number is \[ N=\frac{7\cdot 10^5-35}{49}=14285 \]